OpenStudy (anonymous):

The sum of a number and its reciprocal is 8. Find this number.

OpenStudy (anonymous):

are there choices

OpenStudy (anonymous):

x +1/x = 8 1/x = 8 - x 1 = 8x - x^2 x^2 - 8x + 1 = 0 Then use the quadratic formula...

OpenStudy (anonymous):

[1 ± sqrt(33)] / 2 [-1 ± sqrt(33)] / 2 [-8 ± sqrt(60)] / 2 [8 ± sqrt(60)] / 2

OpenStudy (anonymous):

well since its -b as the first part of quadratics and it turns to be positive 8 the answer is the last one...